### v1

# Graphical sequences and plane trees

Abstract.
A sequence $d_{1}\leqslant\cdots\leqslant d_{n}$ is graphical if it is the degree sequence of a graph. Balister, the second author, Groenland, Johnston and Scott showed that there are asymptotically $C4^{n}/n^{3/4}$ such sequences. However, the constant $C$ involves a probability that is only approximated. Using random walks and limit theory for infinitely divisible probability distributions, we describe $C$ in terms of Walkup’s formula for the number of rooted, unlabelled and cyclically distinct plane trees.
Key words and phrases:
asymptotic enumeration; infinite divisibility; graphical sequence; integrated random walk; lattice point; majorization; persistence probability; plane tree; random walk; renewal sequence
2010 Mathematics Subject Classification:
05A15; 05A16; 05A17; 05C30; 11P21; 51M20; 52B05; 60E07; 60G50; 60K05
1. Introduction
A sequence $d_{1}\leqslant\cdots\leqslant d_{n}$ is graphical if there is a graph on $n$ vertices with this degree sequence. Conditions under which sequences are graphical are well-known; see Havel [11], Hakimi [10] and Erdős and Gallai [8]. We let $\mathcal{G}_{n}$ denote the number of such sequences.
Balister, the second author, Groenland, Johnston and Scott [2] showed that $n^{3/4}\mathcal{G}_{n}/4^{n}\to C$. The constant $C$, however, is expressed in terms of a probability $\rho$ that is estimated, but not calculated; see (1.2) below.
We show that the asymptotics of $\mathcal{G}_{n}$ can be expressed more explicitly, in terms of the number ${\mathcal{T}}_{n}$ of unlabelled, rooted plane trees with $n$ edges, viewed up to cyclical shifts of subtrees around the root, as in Figure 1.
Theorem 1.
As $n\to\infty$, the number $\mathcal{G}_{n}$ of graphical sequences of length $n$ satisfies
$\frac{n^{3/4}}{4^{n}}\mathcal{G}_{n}\to\frac{\Gamma(3/4)}{2^{5/2}\pi}\exp\left(\sum_{k=1}^{\infty}\frac{{\mathcal{T}}_{k}}{k4^{k}}\right).$
[…]
For convenience, we put
$\xi=\sum_{k=1}^{\infty}\frac{{\mathcal{T}}_{k}}{k4^{k}}.$
Using Pólya’s enumeration theorem, Walkup [22] showed that
${\mathcal{T}}_{n}=\frac{1}{n}\sum_{d|n}{2d-1\choose d}\phi(n/d),$ (1.1)
where $\phi(n)$ is Euler’s totient function.
By von Sterneck’s subset counting formulas (see, e.g., [1, 17]), ${\mathcal{T}}_{n}$ is also the number of submultisets of $\{0,1,\ldots,n-1\}$ of size $n$ that sum to 00 mod $n$. This fact plays a key role in our arguments, as it allows us to connect plane trees with lattice walks and random walk bridges.
Figure 1. The ${\mathcal{T}}_{3}=4$ rooted, unlabelled and cyclically distinct plane trees with 3 edges.
There are other equivalent combinatorial descriptions of ${\mathcal{T}}_{n}$, related to necklaces, set partitions, circulant matrices, Hall’s theorem (see Brualdi and Newman [3]), etc. See, e.g., the OEIS entry [19].
Using the values for ${\mathcal{T}}_{n}$ given in [19], it follows that
$\frac{n^{3/4}}{4^{n}}\mathcal{G}_{n}\to C\approx 0.09910,$
which is consistent with the non-rigorous estimate reported in [2].
1.1. Random walks
To our knowledge, Erdős and Moser were the first to observe a connection between random walks and degree sequences, in the context of graph tournaments, as discussed in Moon [15].
A similar connection with random walks is used in [2], to show that
$\frac{n^{3/4}}{4^{n}}\mathcal{G}_{n}\to\frac{\Gamma(3/4)}{2^{5/2}\pi}\frac{1}{\sqrt{1-\rho}},$ (1.2)
where $\rho$ is defined as follows.
Let $Y_{n}$ be a lazy simple random walk, started at $Y_{0}=0$, with increments $Y_{k+1}-Y_{k}$ equal to ±$\pm 1$ with probability $1/4$, and 00 with probability $1/2$. Let
$\tau=\inf\{k\geqslant 1:Y_{k}=0,\>A_{k}\leqslant 0\},$
where $A_{k}=\sum_{i=1}^{k}Y_{i}$ is the area after $k$ steps. Then
$\rho={\mathbb{P}}(A_{\tau}=0).$ (1.3)
In proving Theorem 1, we obtain the following description of $\rho$.
Theorem 2.
We have that $\rho=1-\exp(-2\xi)$.
1.2. Ordered graphical sequences
The polytope $D_{n}$ of degree sequences was introduced by Koren [14], and studied by Peled and Srinivasan [16].
Stanley [21, Corollary 3.4] found a formula for the number of lattice points in $D_{n}\cap\mathbb{Z}^{n}$ with even sum, expressed in terms of quasi-forests. Such points are associated with ordered graphical sequences $d_{1},\ldots,d_{n}$.
On the other hand, Theorem 1 relates the number of unordered graphical sequences $d_{1}\leqslant\cdots\leqslant d_{n}$ to plane trees.
1.3. Acknowledgments
MB is supported by a Clarendon Fund Scholarship. SD acknowledges the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen). BK is supported by a Florence Nightingale Bicentennial Fellowship (Oxford Statistics) and a Senior Demyship (Magdalen College).
2. Infinite divisibility
A random variable $X$ is infinitely divisible (see, e.g., [9]) if, for all $n\geqslant 1$, there are independent and identically distributed $X_{i}$ such that $\sum_{i=1}^{n}X_{i}$ and $X$ are equal in distribution.
Our proof of Theorem 1 is based on limit theory for discrete, infinitely divisible probability distributions. This theory was developed by Hawkes and Jenkins [12] (cf. Embrechts and Hawkes [7]).
Suppose that a positive sequence $(1=a_{0},a_{1},\ldots)$ is summable, and so proportional to probability distribution $\pi(n)$ on the non-negative integers $n\geqslant 0$. Then $\pi$ is infinitely divisible if and only if $(a_{0},a_{1},\ldots)$ has a non-negative log transform $(a^{*}_{1},a^{*}_{2},\ldots)$, such that
$\sum_{n=0}^{\infty}a_{n}x^{n}=\exp\left(\sum_{k=1}^{\infty}\frac{a^{*}_{k}}{k}x^{k}\right).$ (2.1)
As discussed in [7], this is special case of the Lévy–Khintchine representation. In terms of the associated generating functions $A(x)=\sum_{n}a_{n}x^{n}$ and $A^{*}(x)=\sum_{n}a_{n}^{*}x^{n}$, (2.1) says that
$A^{*}(x)=x\frac{d}{dx}\log A(x),$
or, equivalently, we have the recurrence
$na_{n}=\sum_{i=1}^{n}a^{*}_{i}a_{n-i}.$
The following limit theorem is proved in [12].
A positive sequence $\vartheta(n)$ is regularly varying, with index $\gamma$, if
$\lim_{n\to\infty}\frac{\vartheta(\lfloor xn\rfloor)}{x^{\gamma}\vartheta(n)}=1,\quad\quad\forall x>0.$
Lemma 3 ([12, Theorem 3.1]).
Suppose that $a_{n}^{*}$ is regularly varying, with index $\gamma<0$. Then, as $n\to\infty$,
$\frac{a^{*}_{n}}{na_{n}}\to\exp\left(-\sum_{k=1}^{\infty}\frac{a^{*}_{k}}{k}\right).$
This result extends the work of Wright [23, 24, 25]. The proof is based on the analysis of Chover, Ney and Wainger [4].
To prove Theorem 1, we find $a_{n}$, so that $\rho$ is the limit in Lemma 3. Specifically, $a_{n}=\mathcal{B}_{n}/4^{n}$, where $\mathcal{B}_{n}$ is the number of graphical bridges of length $2n$, as defined in Section 3 below. Then, as we will see, $a^{*}_{n}=2{\mathcal{T}}_{n}/4^{n}$, leading to the appearance of ${\mathcal{T}}_{n}$ in Theorem 1.
3. Graphical bridges
Suppose that $X=(X_{0},\ldots,X_{2n})$ is a walk with ±$\pm 1$ increments.
Definition 4.
We call
$\sigma(X)=\frac{1}{2}\sum_{i=1}^{n}X_{2i}$
the diamond area of $X$.
In fact, $\sigma(X)$ is the usual area of the lazy version $\Lambda(X)$ of $X$, defined as follows. Let $\Delta_{i}=X_{i}-X_{i-1}$, for $i\geqslant 1$, be the increments of $X$. The increments $\Delta^{\prime}_{i}$ of $\Lambda(X)$ are the averages $\Delta^{\prime}_{i}=(\Delta_{2i}+\Delta_{2i-1})/2$, for $i\geqslant 1$. Note that, if $X$ has increments ±$\pm 1$ with probability $1/2$, then $\Lambda(X)$ has increments ±$\pm 1$ with probability $1/4$, and 00 with probability $1/2$.
The reason for the name is that $\sigma(X)$ is the signed area under the walk, that lies outside a certain pattern of diamonds (a rotated checkerboard), as depicted in Figure 2.
Figure 2. A bridge $B$ with diamond area $\sigma(B)=2$, and its lazy version $\Lambda(B)$ below.
For walks $X=(X_{0},X_{1},\ldots)$, we let $X^{(k)}=(X_{0},X_{1},\ldots,X_{k})$.
Definition 5.
A bridge $B=(B_{0},\ldots,B_{2n})$ of length $2n$ with increments ±$\pm 1$ is graphical if $\sigma(B)=0$ and $\sigma_{2k}=\sigma(B^{(2k)})\geqslant 0$, for all $1\leqslant k<n$. A graphical bridge $B$ is irreducible if $\sigma_{2k}>0$, for all $1\leqslant k<n$. We let $\mathcal{B}_{n}$ be the number of graphical bridges of length $2n$. See Figure 3.
Figure 3. There are $\mathcal{B}_{5}=38$ graphical bridges of length $10$. Of these, 32 stay within the string of grey diamonds. The other 6 are depicted above. The top two bridges are irreducible. All others have two irreducible parts.
3.1. Relation to $\rho$
Recall that $\rho$ is defined in terms of a lazy random walk. We can rephrase $\rho$ in terms of a standard, simple random walk $(X_{n})$ with ±$\pm 1$ increments, started from $X_{0}=0$, using its diamond area. Specifically, let $\sigma_{2k}=\sigma(X^{(2k)})$ denote the diamond area accumulated after $2k$ steps. Since, as discussed above, the diamond area of $X$ is the usual area of its lazy version $\Lambda(X)$, it follows that
$\rho={\mathbb{P}}(\sigma_{2\kappa}=0),\quad\quad\quad\kappa=\inf\{k\geqslant 1:X_{2k}=0,\>\sigma_{2k}\leqslant 0\}.$
To see how $\rho$ relates to graphical bridges, observe that a bridge $B=(B_{0},\ldots,B_{2n})$ (with ±$\pm 1$ increments) is graphical if and only if $\sigma(B)=0$ and $\sigma_{2k}=\sigma(B^{(2k)})\geqslant 0$, for all times $2k$ for which $B_{2k}=0$, since $\sigma$ is monotone between such diagnostic times. Therefore, roughly speaking, $\rho$ is the probability that, the first time a random walk trajectory is in danger of not satisfying $\sigma_{2k}\geqslant 0$, at such a diagnostic time $2k$, the condition is in fact satisfied with equality $\sigma_{2k}=0$.
Consider now a random walk $X=(X_{0},X_{1},\ldots)$. If $X_{2k}=\sigma_{2k}=0$ at the first diagnostic time $2k<2n$, then $(X_{0},\ldots,X_{2n})$ is a graphical bridge if and only if $(X_{2k},\ldots,X_{2n})$ is a graphical bridge. Therefore, times at which $X_{2k}=\sigma_{2k}=0$ are of crucial importance, as they decompose a graphical bridge into its irreducible parts, and correspond to renewal times in the analysis.
Definition 6.
We let ${\mathcal{I}}_{n}$ denote the number of irreducible parts in a uniformly random graphical bridge of length $2n$.
Our arguments in [6, Proposition 21] can be adapted to show that
[…]
where ${\mathcal{X}}$ is a negative binomial with parameters $r=2$ and $p=1-\rho$. The reason is that, with high probability, renewal times occur only very close to the start and end of the bridge. Moreover, these numbers are approximately independent and geometric.
3.2. Calculating $\rho$
As we will see, the probability $\rho$ is related to the asymptotics of $\mathcal{B}_{n}$, which we analyze using Lemma 3.
As we have discussed, a graphical bridge can be decomposed into a series of irreducible parts. In other words, $\mathcal{B}_{n}$ is a renewal sequence (see, e.g., [9]). Let $\mathcal{B}_{n}^{(m)}$ be the number of graphical bridges of length $2n$ with exactly $m$ irreducible parts. Then, letting $\mathcal{B}(x)=\sum_{n}\mathcal{B}_{n}x^{n}$ and $\mathcal{B}^{(m)}(x)=\sum_{n}\mathcal{B}_{n}^{(m)}x^{n}$ denote the associated generating functions, we have that
$\mathcal{B}(x)=\frac{1}{1-\mathcal{B}^{(1)}(x)}.$
More specifically,
$\mathcal{B}_{n}^{(m)}=[x^{n}](\mathcal{B}^{(1)}(x))^{m}$
is the number of ways to concatenate a series of $m$ irreducible graphical bridges into a graphical bridge with total length $2n$.
Renewal sequences $a_{n}$ are infinitely divisible. In this case, the log transform $a_{n}^{*}$ has a special form (see, e.g., [12, 13]). In our current setting,
$\frac{\mathcal{B}_{n}^{*}}{n\mathcal{B}_{n}}=\sum_{m}\frac{1}{m}\frac{\mathcal{B}_{n}^{(m)}}{\mathcal{B}_{n}}={\mathbb{E}}[1/{\mathcal{I}}_{n}].$ (3.2)
Therefore, by (3.1) and (3.2), we find that
$\frac{\mathcal{B}^{*}_{n}}{n\mathcal{B}_{n}}\to 1-\rho.$ (3.3)
In proving Theorems 1 and 2, the following is key.
Proposition 7.
The log transform of $\mathcal{B}_{n}$ is $\mathcal{B}^{*}_{n}=2{\mathcal{T}}_{n}$.
This result is useful, since ${\mathcal{T}}_{n}$ has formula (1.1), and simple asymptotics. We note that $2{\mathcal{T}}_{n}$ has an OEIS entry [20].
Table 1. Initial values of $\mathcal{B}_{n}$ and its log transform $\mathcal{B}^{*}=2{\mathcal{T}}_{n}$.
$n$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$ $9$
$\mathcal{B}_{n}$ $1$ $2$ $4$ $8$ $17$ $38$ $92$ $236$ $643$ $1834$
$2{\mathcal{T}}_{n}$ 00 $2$ $4$ $8$ $20$ $52$ $160$ $492$ $1620$ $5408$
The proof of Proposition 7 is postponed until Section 5, but let us note here that our main results follow.
Proof of Theorems 1 and 2.
By (1.1) and Proposition 7,
$\frac{\mathcal{B}^{*}_{n}}{4^{n}}\sim\frac{1}{n4^{n}}{2n\choose n}\sim\frac{1}{\sqrt{\pi}n^{3/2}}$
is regularly varying with index $\gamma=-3/2$. By Lemma 3, it follows that
$\frac{\mathcal{B}^{*}_{n}}{n\mathcal{B}_{n}}\to e^{-2\xi}.$ (3.4)
Combining (3.3) and (3.4), we obtain Theorem 2. Substituting this into (1.2), our main result Theorem 1 follows. ∎
4. Combinatorial lemmas
In this section, we prove two combinatorial lemmas, showing that $2{\mathcal{T}}_{n}$ can be described in terms of areas $\alpha$ below lattice paths $L$ and diamond areas $\sigma$ under bridges $B$.
4.1. Lattice paths
Suppose that $L$ is an $\uparrow$,$\rightarrow$ , → lattice walk from $(a_{1},b_{1})$ to $(a_{2},b_{2})$, for some integers $a_{1}\leqslant a_{2}$ and $b_{1}\leqslant b_{2}$. We let $\alpha(L)$ denote the area of the region between $L$ and the lines $x=a_{2}$ and $y=b_{1}$.
Definition 8.
We let $\mathcal{N}_{n}$ be the number of lattice paths $L$ from $(0,0)$ to $(n,n)$ such that $\alpha(L)\equiv 0$ mod $n$.
To see the relationship between $\mathcal{N}_{n}$ and ${\mathcal{T}}_{n}$, first note that the area of a lattice path $L$ from $(0,0)$ to $(n,n)$ is
$\alpha(L)=\sum_{i}u_{i},$ (4.1)
summing over the sequence $u_{1}\leqslant\cdots\leqslant u_{n}$, where $u_{i}$ is the number of $\uparrow$ steps before the $i$th $\rightarrow$ step of $L$. In other words, if we picture $L$ as a bar graph, then the $u_{i}$ are the heights of the bars. Hence, $\mathcal{N}_{n}$ is the number of submultisets of $\{0,1,\ldots,n\}$ of size $n$ that sum to 00 mod $n$. On the other hand, recall that ${\mathcal{T}}_{n}$ is equal to the number of submultisets $\{0,1,\ldots,n-1\}$ of size $n$ that sum to 00 mod $n$.
Using these observation, we show the following.
Lemma 9.
We have that $\mathcal{N}_{n}=2{\mathcal{T}}_{n}$.
Proof.
One way to prove this is to use the fact (see, e.g., [1, 17]) that
$\mathcal{M}_{n,k}=\frac{1}{n}\sum_{d|k,n}{(n+k)/d-1\choose k/d}\phi(d)$
is the number of submultisets of $\{0,1,\ldots,n-1\}$ of size $k$ that sum to 00 mod $n$. Therefore, using (1.1), and considering the number $k$ of $u_{i}<n$,
$\mathcal{N}_{n}=\sum_{k=0}^{n}\mathcal{M}_{n,k}={\mathcal{T}}_{n}+\sum_{k=1}^{n}\mathcal{M}_{n,k}=2{\mathcal{T}}_{n},$
where the last equality follows after reversing the order of summation.
We also give the following combinatorial proof. Similar reasoning as (4.1) shows that ${\mathcal{T}}_{n}$ is equal to the number of lattice paths $L$ from $(0,0)$ to $(n,n-1)$ with $\alpha(L)\equiv 0$ mod $n$. Indeed, submultisets $10\leqslant u_{1}\leqslant\cdots\leqslant u_{n}\leqslant n-$ 1 correspond to such $L$ with $\rightarrow$ steps at heights $u_{i}$, and so $\alpha(L)=\sum_{i}u_{i}\equiv 0$ mod $n$.
Let $\mathcal{L}_{n}^{\uparrow}$ (resp. $\mathcal{L}_{n}^{\rightarrow}$) be the number of lattice paths $L$ from $(0,0)$ to $(n,n)$ such that $\alpha(L)\equiv 0$ mod $n$, ending with an $\uparrow$ (resp. $\rightarrow$) step. Then $\mathcal{N}_{n}=\mathcal{L}_{n}^{\uparrow}+\mathcal{L}_{n}^{\rightarrow}$. To conclude, we show that $\mathcal{L}_{n}^{\uparrow}=\mathcal{L}_{n}^{\rightarrow}=\mathcal{T}_{n}$. Indeed, $\mathcal{L}_{n}^{\uparrow}=\mathcal{L}_{n}^{\rightarrow}$ follows by symmetry, reflecting over $x=y$ Finally, $\mathcal{L}_{n}^{\uparrow}=\mathcal{T}_{n}$, since if $L$ from $(0,0)$ to $(n,n)$ has last step $\uparrow$, then removing this step yields a corresponding $L^{\prime}$ from $(0,0)$ to $(n,n-1)$ with the same area as $L$. ∎
4.2. Bridges
Next, we connect $\mathcal{N}_{n}$ to bridges.
Definition 10.
We let $\mathcal{N}_{n}^{\prime}$ denote the number of bridges $B$ of length $2n$, with diamond area $\sigma(B)\equiv 0$ mod $n$.
Lemma 11.
We have that $\mathcal{N}_{n}=\mathcal{N}_{n}^{\prime}$.
Proof.
To obtain a correspondence between bridges $B=(B_{0},\ldots,B_{2n})$ of length $2n$ such that $\sigma(B)\equiv 0$ mod $n$ and lattice paths $L$ from $(0,0)$ to $(n,n)$ with $a(L)\equiv 0$ mod $n$, we proceed as follows.
First, let $\Delta_{i}=B_{i}-B_{i-1}$ be the $i$th increment of $B$. For $1\leqslant k\leqslant n$, put
$B^{\mathrm{odd}}_{k}=\sum_{i=1}^{k}\Delta_{2i-1},\quad\quad B^{\mathrm{even}}_{k}=\sum_{i=1}^{k}\Delta_{2i},$
so that $B^{\mathrm{odd}}=(B^{\mathrm{odd}}_{1},\ldots,B^{\mathrm{odd}}_{n})$ and $B^{\mathrm{even}}=(B^{\mathrm{even}}_{1},\ldots,B^{\mathrm{even}}_{n})$ are the walks with the odd and even increments of $B$, respectively.
Next, rotate $B^{\mathrm{odd}}$ counterclockwise by $\pi/2$ to obtain a lattice path $L_{1}$ from $(0,0)$ to $(n-\ell,\ell)$ for some $\ell$ (since $B^{\mathrm{odd}}$ has $n$ steps). Likewise, rotate $B^{\mathrm{even}}$ counterclockwise by $\pi/2$ to obtain a lattice walk $L_{2}$ from $(n-\ell,\ell)$ to $(n,n)$ (since $B$ is a bridge). Let $L$ be the concatenation of $L_{1}$ and $L_{2}$. This procedure is depicted in Figure 4.
Geometric considerations (see Figure 5) imply that
$\alpha(L)=\alpha(L_{1})+\alpha(L_{2})+\ell^{2}.$
Furthermore, observe that, by definition,
$\sigma(B)=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\left(\Delta_{2j-1}+\Delta_{2j}\right)=\frac{1}{2}\sum_{k=1}^{n}(B^{\mathrm{odd}}_{k}+B^{\mathrm{even}}_{k}),$
which equals the signed area enclosed by $B^{\mathrm{odd}}$ and $-B^{\mathrm{even}}$, as depicted in Figure 5. Therefore, it can be seen that
$\sigma(B)=\alpha(L_{1})-[\ell(n-\ell)-\alpha(L_{2})]=\alpha(L)-\ell n.$
Since $\alpha(L)\equiv\sigma(B)$ mod $n$, this completes the proof. ∎
Figure 4. The bijection in Lemma 11.
Figure 5. Calculating areas in Lemma 11.
5. The log transform
We are now in a position to prove our key result Proposition 7, which identifies $\mathcal{B}_{n}^{*}=2{\mathcal{T}}_{n}$ as the log transform of $\mathcal{B}_{n}$.
Proof of Proposition 7.
The first step is to observe (see [13, §3.3]) that the ratio $\mathcal{B}_{n}^{*}/n\mathcal{B}_{n}$ in (3.3) and (3.4) can be calculated as follows. Consider a uniformly random graphical bridge $B$ of length $2n$, and let $\ell$ be an independent, uniformly random integer in $\{1,2,\ldots,n\}$. Then $\mathcal{B}^{*}_{n}/n\mathcal{B}_{n}$ is the probability that the first irreducible part in $B$ is of length at most $2\ell$.
Therefore, $\mathcal{B}_{n}^{*}$ is the number of ordered pairs $(B,i)$, where $B$ is a graphical bridge of length $2n$, whose first irreducible part is of length $2j$, and $0\leqslant i<j$.
For each such pair $(B,i)$, we let $\phi(B,i)$ be the bridge obtained from $B$, via a cyclical shift to the right by $2i$. We claim that $\phi$ is a bijection from the set of pairs $(B,i)$ to the set of bridges with diamond areas $\sigma\equiv 0$ mod $n$. By Lemmas 9 and 11, this will complete the proof.
The mapping $\phi$ is injective, since we do not shift past the first irreducible part in a $B$. To see that $\phi$ is surjective, first note that, as we shift the $x$-axis up or down by multiples of 2, the diamond area $\sigma$ changes by multiples of $n$. Therefore, if $B$ has $\sigma\equiv 0$ mod $n$, then some cyclical shift of $B$ is graphical. To see this, we select the shift that sets $\sigma=0$, and then apply Raney’s lemma [18], as in Figure 6. We omit the details, as they are similar to the proof of [5, Lemma 8]. ∎
Figure 6. A graphical bridge $B$ (top left) of length 10, with first irreducible part of length 8. The bridge $B=\phi(B,0)$. Its other shifts $\phi(B,i)$, for $41\leqslant i<$ 4 are also depicted. All bridges have diamond area divisible by 5. To find the inverse mapping $\phi^{-1}$, we shift the $x$-axis, by some factor of 2, so that the diamond area is equal to 0, and then apply Raney’s lemma.