### v2

# Graphical sequences and plane trees

Abstract.
Balister, the second author, Groenland, Johnston and Scott recently showed that there are asymptotically $C4^{n}/n^{3/4}$ many unordered sequences that occur as degree sequences of graphs. Combining limit theory for infinitely divisible distributions with a new bijective connection between a class of random walk trajectories and a subset counting formula from additive number theory, we describe $C$ in terms of Walkup’s number of rooted plane trees. The bijection is related to an instance of the Lévy–Khintchine formula. Our main result complements a result of Stanley, that ordered graphical sequences are related to quasi-forests.
Key words and phrases:
asymptotic enumeration; infinite divisibility; graphical sequence; integrated random walk; lattice point; majorization; persistence probability; plane tree; random walk; renewal sequence
2010 Mathematics Subject Classification:
05A15; 05A16; 05A17; 05C30; 11P21; 51M20; 52B05; 60E07; 60G50; 60K05
1. Introduction
In this work, we reveal a connection between trees embedded in the plane and the degree sequences of graphs. The proof involves an interplay between additive number theory, random walks, renewal theory and infinite divisibility. In particular, we use the Lévy–Khintchine formula, from the theory of Lévy processes, to obtain a result in the area of combinatorial asymptotic enumeration. We expect this strategy to find more applications.
A sequence $d_{1}\leqslant\cdots\leqslant d_{n}$ is graphical if there is a graph on $n$ vertices with this degree sequence. Conditions under which sequences are graphical are well-known; see Havel [11], Hakimi [10] and Erdős and Gallai [7]. Recently, Balister, the second author, Groenland, Johnston and Scott [2] showed that the number $\mathcal{G}_{n}$ of such sequences satisfies $n^{3/4}\mathcal{G}_{n}/4^{n}\to C$. The constant $C$ is expressed as a certain random walk probability $\rho$; see (1.3) below.
We show that $C$ can be expressed in terms of the number ${\mathcal{T}}_{n}$ of rooted unlabeled (cyclically distinct) plane trees with $n$ edges; see Figure 1 below.
Theorem 1.
As $n\to\infty$, we have that
$\frac{n^{3/4}}{4^{n-1}}\mathcal{G}_{n}\to\frac{e^{\omega}}{\Gamma(1/4)},$
where
$\omega=\sum_{k=1}^{\infty}\frac{{\mathcal{T}}_{k}}{k4^{k}}.$ (1.1)
[…]
Using Pólya’s enumeration theorem, Walkup [19] showed that
${\mathcal{T}}_{n}=\frac{1}{n}\sum_{d|n}{2d-1\choose d}\phi(n/d),$ (1.2)
where $\phi(n)$ is Euler’s totient function. Thus, we call $\omega$ Walkup’s constant.
By von Sterneck’s subset counting formulas (see, e.g., [1, 16]), ${\mathcal{T}}_{n}$ is also the number of submultisets of $\{0,1,\ldots,n-1\}$ of size $n$ that sum to 00 mod $n$. This is key in our arguments, since, as we will see, it can be used to connect plane trees with lattice paths.
Figure 1. The ${\mathcal{T}}_{3}=4$ rooted unlabeled plane trees with 3 edges.
1.1. Degree sequences, via random walks
Erdős and Moser observed a connection between random walks and degree sequences, in the context of graph tournaments, as discussed in Moon [14].
A similar connection with random walks is used in [2] to show that
$\frac{n^{3/4}}{4^{n}}\mathcal{G}_{n}\to\frac{\Gamma(3/4)}{2^{5/2}\pi}\frac{1}{\sqrt{1-\rho}},$ (1.3)
where $\rho$ is defined as follows.
Let $Y_{n}$ be a lazy simple symmetric random walk, started at $Y_{0}=0$, with increments $Y_{k+1}-Y_{k}$ equal to ±$\pm 1$ with probability $1/4$, and 00 with probability $1/2$. Let
$\tau=\inf\{k\geqslant 1:Y_{k}=0,\>A_{k}\leqslant 0\},$
where $A_{k}=\sum_{i=1}^{k}Y_{i}$ is the area after $k$ steps. Then
$\rho={\mathbb{P}}(A_{\tau}=0).$ (1.4)
In proving Theorem 1, we obtain the following description of $\rho$, in terms of Walkup’s constant $\omega$.
Theorem 2.
We have that $\rho=1-e^{-2\omega}$.
We note that Theorem 1 follows by combining Theorem 2 with (1.3), and noting that $\Gamma(3/4)=\pi\sqrt{2}/\Gamma(1/4)$.
1.2. Proof overview
There is a natural representation of the lazy simple symmetric random walk $Y_{n}$, discussed above, in terms of a simple symmetric random walk $X_{n}$ of twice the length. In this representation, the area process $A_{k}=\sum_{i=1}^{k}Y_{i}=\frac{1}{2}\sum_{i=1}^{k}X_{2i}$. This relationship has a geometric interpretation; see the diamond area discussed in Section 3 below.
If $X_{2n}=0$ and $A_{k}\geqslant 0$ at all times $0\leqslant k\leqslant n$ for which $X_{2k}=0$, we call $\{X_{k},0\leqslant k\leqslant 2n\}$ a graphical walk of length $2n$. A graphical bridge is a graphical walk whose area returns to 00 at the end and a graphical meander is a graphical walk whose area never returns to 00. Times $k$ when $X_{2k}=A_{k}=0$ play the role of renewal times, since initially $X_{0}=A_{0}=0$. Hence a graphical walk can be decomposed into a series of graphical bridges and one final graphical meander. That is, ${\mathcal{W}}_{n}$ is a delayed renewal sequence, where the length of the final meander can be understood as the “delay.” The term $\Gamma(3/4)/2^{5/2}\pi$ in (1.3) is related to the number $\mathcal{M}_{n}$ of graphical meanders.
To describe the other term $1/\sqrt{1-\rho}$ in (1.3), we analyze the renewal sequence (without a delay) $\mathcal{B}_{n}$ of graphical bridges. We use the asymptotic transference theory, developed by Hawkes and Jenkins [12] (cf. Embrechts and Hawkes [6]), for infinitely divisible distributions. We call this method the Lévy–Khintchine method (see Section 2 below), as it is related to the Lévy–Khintchine transform, which associates each such distribution with a corresponding Lévy process, via another measure called its Lévy measure.
We show that the Lévy–Khintchine transform $\mathcal{B}_{n}^{*}$ of $\mathcal{B}_{n}$ is related to ${\mathcal{T}}_{n}$. In fact, $\mathcal{B}_{n}^{*}=2{\mathcal{T}}_{n}$. More specifically, we find that $p_{n}=e^{-2\omega}\mathcal{B}^{n}/4^{n}$ is an infinitely divisible probability distribution with Lévy measure $\nu_{n}=2{\mathcal{T}}_{n}/n4^{n}$. The measure $\nu_{n}$ is regularly varying, and it then follows by [12] that $p_{n}\sim\nu_{n}$.
The proof that $\mathcal{B}_{n}^{*}=2{\mathcal{T}}_{n}$ involves a characterization (see Lemma 3 below) of the Lévy–Khintchine transform of renewal sequences and some delicate combinatorics, illustrated in Figures 4 and 5 below. Roughly speaking, by interleaving the steps of certain lattice paths, we relate the set of cyclical shifts of graphical bridges and the set of Walkup’s plane trees, via von Sterneck’s formula. Our arguments are geometric, using the diamond area mentioned above. See Figures 2, 3 and 6.
Since $\mathcal{B}_{n}$ is a renewal sequence, $\nu_{n}$ is also the harmonic renewal measure associated with the sequence $\mathcal{B}_{n}^{(1)}$ of irreducible graphical bridges. Using this, it follows that $e^{-2\omega}\nu_{n}/p_{n}$ is the harmonic moment ${\mathbb{E}}[1/{\mathcal{I}}_{n}]$ of the number ${\mathcal{I}}_{n}$ of irreducible graphical bridges in a uniformly random graphical bridge of length $n$ and therefore ${\mathbb{E}}[1/{\mathcal{I}}_{n}]\to e^{-2\omega}$ (since $p_{n}\sim\nu_{n}$).
Moreover, it can be shown that ${\mathcal{I}}_{n}$ converges to a negative binomial with parameters $r=2$ and $p=1-\rho$. Intuitively, renewal times will only occur near the start and end of the bridge. The numbers of such times on either side of the bridge are approximately independent and geometric with $p=1-\rho$. Therefore, ${\mathbb{E}}[1/{\mathcal{I}}_{n}]\to 1-\rho$, and so we find that $1-\rho=e^{-2\omega}$, yielding Theorem 2. Hence, $1/\sqrt{1-\rho}=e^{\omega}$, and so Theorem 1 follows by (1.3).
1.3. Ordered graphical sequences
The polytope $D_{n}$ of degree sequences was introduced by Koren [13], and studied by Peled and Srinivasan [15]. Stanley [18, Corollary 3.4] found a formula for the number of lattice points in $D_{n}\cap\mathbb{Z}^{n}$ with even sum, expressed in terms of quasi-forests. Such points are associated with ordered graphical sequences $d_{1},\ldots,d_{n}$. On the other hand, Theorem 1 relates the number of unordered graphical sequences $d_{1}\leqslant\cdots\leqslant d_{n}$ and Walkup’s plane trees.
1.4. Numerics
Finally, let us note that the expression for $C$ in Theorem 1 can be used to approximate $\mathcal{G}_{n}$, for large $n$. Observing that the formula (1.2) for ${\mathcal{T}}_{n}$ is dominated by the $d=n$ term, one can show that
[…]
See the discussion of this work in [2] for more details.
1.5. Acknowledgments
MB is supported by a Clarendon Fund Scholarship. SD acknowledges the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen). BK was supported by a Florence Nightingale Bicentennial Fellowship (Oxford Statistics) and a Senior Demyship (Magdalen College).
2. The Lévy–Khintchine method
A random variable $X$ is infinitely divisible if, for all $n\geqslant 1$, there are independent and identically distributed $X_{i}$ such that $\sum_{i=1}^{n}X_{i}$ and $X$ are equal in distribution (see, e.g.,[8]).
Suppose that a positive sequence $(1=a_{0},a_{1},\ldots)$ is summable, so that it is proportional to a probability distribution $\pi_{n}$ on the non-negative integers $n\geqslant 0$. As is well-known (see, e.g., [12]) such a $\pi_{n}$ is infinitely divisible if and only if, for some non-negative sequence $(a^{*}_{1},a^{*}_{2},\ldots)$, we have that
$\sum_{n=0}^{\infty}a_{n}x^{n}=\exp\left(\sum_{k=1}^{\infty}\frac{a^{*}_{k}}{k}x^{k}\right).$ (2.1)
In this case, we call $(a_{0},a_{1},\dots)$ an infinitely divisible sequence. By taking derivatives with respect to $x$ on both sides of (2.1) and comparing coefficients, we see that (2.1) is equivalent to the recursion
$na_{n}=\sum_{i=1}^{n}a^{*}_{i}a_{n-i},\quad\quad n\geqslant 1.$ (2.2)
Since, as discussed in [6], (2.1) is a special case of the Lévy–Khintchine formula, we call $a_{n}^{*}$ the Lévy–Khintchine transform of $a_{n}$. We note that, in fact, $\nu_{n}=a^{*}_{n}/n$ is the Lévy measure (see, e.g., [4]) associated with the Lévy process $\{{\mathcal{X}}_{t}:t\geqslant 0\}$, for which the law of the process ${\mathcal{X}}_{1}$ at time $t=1$ is described by the infinitely divisible probability measure
$\pi_{n}=a_{n}\exp\left(-\sum_{k=1}^{\infty}\frac{a^{*}_{k}}{k}\right),\quad\quad n\geqslant 0.$
In combinatorial settings, it is convenient to consider $a_{n}=A_{n}/\alpha^{n}$, where $A_{n}$ enumerates a class of objects of size $n$ that has exponential growth rate $\alpha$. If $a_{n}$ is infinitely divisible, then $A_{n}$ and $A^{*}_{n}=\alpha^{n}a_{n}^{*}$ also satisfy (2.1) and (2.2), and naturally we also call $A^{*}_{n}$ the Lévy–Khintchine transform of $A_{n}$.
A positive sequence $\vartheta(n)$ is regularly varying, with index $\gamma$, if
$\lim_{n\to\infty}\frac{\vartheta(\lfloor xn\rfloor)}{\vartheta(n)}=x^{\gamma},\quad\quad\forall x>0.$
One of our main tools for studying the asymptotics of sequences with a Lévy–Khintchine transform is a result by Hawkes and Jenkins [12], which shows that, if $a_{n}^{*}$ is regularly varying with some index $\gamma<0$, then
$a_{n}\sim\frac{a_{n}^{*}}{n}\exp\left(\sum_{k=1}^{\infty}\frac{a^{*}_{k}}{k}\right).$ (2.3)
In other words, in terms of the underlying Lévy process, the associated infinitely divisible probability measure $\pi_{n}$ and its corresponding Lévy measure $\nu_{n}$ are asymptotically equivalent, $\pi_{n}\sim\nu_{n}$ as $n\to\infty$.
2.1. Renewal sequences
A subfamily of sequences $A_{n}$ that have a Lévy-Khintchine transform is the class of renewal sequences. Such a sequence arises when $A_{n}$ enumerates structures of length $n$ that can be decomposed into a sequence of irreducible parts. In this case, the generating function $A(x)=\sum_{n\geqslant 1}A_{n}x^{n}$ of $A_{n}$ satisfies
$A(x)=\frac{1}{1-A^{(1)}(x)}$
where $A^{(1)}(x)=\sum_{n\geqslant 1}A^{(1)}_{n}x^{n}$ is the generating function of the irreducible structures (see, e.g., [8]).
In our recent work [3], we show that, when $A_{n}$ is a renewal sequence, $A^{*}_{n}$ takes a special form.
Lemma 3 ([3, Lemma 2]).
Suppose that $A_{n}$ is a renewal sequence. Then:
  1. (1)
    the Lévy–Khintchine transform $A_{n}^{*}$ is the number of pairs $(X,m)$, where $X$ is a structure of length $n$ and $0\leqslant m<\ell$, where $\ell=\ell(X)$ is the length of the first irreducible part of $X$, and
  2. (2)
    we have that
    $\frac{A_{n}^{*}}{nA_{n}}={\mathbb{E}}\left[\frac{1}{{\mathcal{I}}_{n}}\right],$ (2.4)
    where ${\mathcal{I}}_{n}$ is the number of irreducible parts in a uniformly random structure of length $n$.
We note that, in terms of the underlying Lévy process, the formula (2.4) says that, when $A_{n}$ is a renewal sequence with growth rate $\alpha$, the Lévy measure $\nu_{n}=A_{n}^{*}/n\alpha^{n}$ is the harmonic renewal measure (see, e.g., [9]) associated with the measure $\mu_{n}=A_{n}^{(1)}/\alpha^{n}$.
With Lemma 3 in hand, one can use (2.3) to deduce the asymptotic growth of a renewal sequence $A_{n}$ from the asymptotic growth of $A^{*}_{n}$, provided that there exists an $\alpha$ for which $A^{*}_{n}/\alpha^{n}$ is regularly varying with index $\gamma<0$. In our experience, the sequence $A^{*}_{n}$ tends to be more tractable than the renewal sequence $A_{n}$ itself, which lies at the core of our method.
2.2. Delayed renewal sequences
A related family of sequences $A_{n}$ is the class of delayed renewal sequences. Such sequences arise when $A_{n}$ enumerates structures of length $n$ that can be decomposed into exactly one special part (the “delay”) and a sequence of irreducible parts. In this case, the generating function $A$ of $A_{n}$ satisfies
$A(x)=\frac{D(x)}{1-A^{(1)}(x)},$
where $D(x)$ is the generating function of the delay and $A^{(1)}(x)$ is the generating function of the irreducible structures.
Although $A_{n}$ itself may not necessarily admit a Lévy–Khintchine transform, we illustrate in this work that the asymptotic growth of $A_{n}$ can still be understood by studying the delay and its renewal structure separately. We believe that this method will be useful in enumerating many other combinatorial sequences of interest.
2.3. Application to $\mathcal{G}_{n}$
As discussed in Section 1.2, the sequence $\mathcal{G}_{n}$ that enumerates graphical sequences is a delayed renewal sequence. The $1/\Gamma(1/4)$ factor in the asymptotics corresponds to the delay. To complete the proof of Theorem 1, it remains to consider the asymptotics of the renewal sequence $\mathcal{B}_{n}$ of graphical bridges of length $2n$, to which we will apply the Lévy–Khintchine method, as described above.
More specifically, we consider $a_{n}=\mathcal{B}_{n}/4^{n}$, and its Lévy–Khintchine transform $a^{*}_{n}$, and show that the limit in (2.3) equals $1-\rho$. Finally, we use Lemma 3 and a bijective argument to prove that $a^{*}_{n}=2{\mathcal{T}}_{n}/4^{n}$, leading to the appearance of ${\mathcal{T}}_{n}$ and $\omega$ in Theorems 1 and 2.
The two parts of Lemma 3 play an essential role. Lemma 3(1) allows us to approach the asymptotics of $\mathcal{B}_{n}$ combinatorially, and Lemma 3(2) opens the door to probability theory. These approaches meet halfway, so to speak, to give an explicit description of the asymptotics of $\mathcal{G}_{n}$.
3. Graphical bridges
Suppose that $X=(X_{0},\ldots,X_{2n})$ is a walk with ±$\pm 1$ increments.
Definition 4.
We call
$\sigma(X)=\frac{1}{2}\sum_{i=1}^{n}X_{2i}$ (3.1)
the diamond area of $X$.
In fact, $\sigma(X)$ is the usual area of the lazy version $\Lambda(X)$ of $X$, defined as follows. Let $\Delta_{i}=X_{i}-X_{i-1}$, for $i\geqslant 1$, be the increments of $X$. The increments $\Delta^{\prime}_{i}$ of $\Lambda(X)$ are the averages $\Delta^{\prime}_{i}=(\Delta_{2i}+\Delta_{2i-1})/2$, for $i\geqslant 1$. Note that, if $X$ has increments ±$\pm 1$ with probability $1/2$, then $\Lambda(X)$ has increments ±$\pm 1$ with probability $1/4$, and 00 with probability $1/2$.
The reason for the name “diamond area” is that $\sigma(X)$ is the signed area under the walk, that lies outside a rotated checkerboard of diamonds, as in Figure 2.
Figure 2. A bridge $B$ with diamond area $\sigma(B)=2$, and its lazy version $\Lambda(B)$ below.
For walks $X=(X_{0},X_{1},\ldots)$, we let $X^{(k)}=(X_{0},X_{1},\ldots,X_{k})$.
Definition 5.
A bridge $B=(B_{0},\ldots,B_{2n})$ of length $2n$ with increments ±$\pm 1$ is graphical if $\sigma(B)=0$ and $\sigma_{2k}=\sigma(B^{(2k)})\geqslant 0$, for all $1\leqslant k<n$. A graphical bridge $B$ is irreducible if $\sigma_{2k}>0$, for all $1\leqslant k<n$. We let $\mathcal{B}_{n}$ be the number of graphical bridges of length $2n$. See Figure 3.
Figure 3. There are $\mathcal{B}_{5}=38$ graphical bridges of length $10$. Of these, 32 stay within the string of grey diamonds centered along the $x$-axis. The other 6 are depicted above. The top two are irreducible. All others have two irreducible parts. Irreducible parts are separated by solid dots.
3.1. Relation to $\rho$
Recall that $\rho$ is defined, in Section 1.1 above, in terms of a lazy random walk. We can rephrase $\rho$ in terms of a standard, simple symmetric random walk $(X_{n})$ with ±$\pm 1$ increments, started from $X_{0}=0$, using its diamond area. Specifically, let $\sigma_{2k}=\sigma(X^{(2k)})$ denote the diamond area accumulated after $2k$ steps. Since, as discussed above, the diamond area of $X$ is the usual area of its lazy version $\Lambda(X)$, it follows that
$\rho={\mathbb{P}}(\sigma_{2\kappa}=0),\quad\quad\quad\kappa=\inf\{k\geqslant 1:X_{2k}=0,\>\sigma_{2k}\leqslant 0\}.$
To see how $\rho$ relates to our problem, we observe that a bridge $B=(B_{0},\ldots,B_{2n})$ (with ±$\pm 1$ increments) is graphical if and only if $\sigma(B)=0$ and $\sigma_{2k}=\sigma(B^{(2k)})\geqslant 0$, for all times $2k$ for which $B_{2k}=0$, since $\sigma$ is monotone between such diagnostic times. Then, roughly speaking, $\rho$ is the probability that, the first time a random walk trajectory is in danger of not satisfying $\sigma_{2k}\geqslant 0$, at such a diagnostic time $2k$, the condition is in fact satisfied with equality $\sigma_{2k}=0$.
Consider now a random walk $X=(X_{0},X_{1},\ldots)$. If $X_{2k}=\sigma_{2k}=0$ at the first diagnostic time $2k<2n$, then $(X_{0},\ldots,X_{2n})$ is a graphical bridge if and only if $(X_{2k},\ldots,X_{2n})$ is a graphical bridge. Therefore, times at which $X_{2k}=\sigma_{2k}=0$ are of crucial importance, as they decompose a graphical bridge into its irreducible parts, and correspond to renewal times in the analysis.
Definition 6.
We let ${\mathcal{I}}_{n}$ denote the number of irreducible parts in a uniformly random graphical bridge of length $2n$.
By (the proof of) [5, Proposition 21] it follows that
[…]
where ${\mathcal{X}}$ is a negative binomial with parameters $r=2$ and $p=1-\rho$. The appearance of the negative binomial random variable can, intuitively, be explained as follows. With high probability, renewal times occur only very close to the start and end of the bridge. Moreover, the number of renewal times on either side are approximately geometric and independent.
3.2. Calculating $\rho$
We observed that a graphical bridge can be decomposed into a series of irreducible parts, and so $\mathcal{B}_{n}$ is a renewal sequence. As discussed in Section 2, this means that the generating function $\mathcal{B}(x)=\sum_{n}\mathcal{B}_{n}x^{n}$ can be expressed as
$\mathcal{B}(x)=\frac{1}{1-\mathcal{B}^{(1)}(x)},$
where $\mathcal{B}^{(1)}(x)=\sum_{n}\mathcal{B}_{n}^{(1)}x^{n}$ is the generating function for the number $\mathcal{B}_{n}^{(1)}$ of irreducible graphical bridges of length $2n$.
Therefore, by Lemma 3(2), we have
$\frac{\mathcal{B}_{n}^{*}}{n\mathcal{B}_{n}}={\mathbb{E}}\left[\frac{1}{{\mathcal{I}}_{n}}\right].$ (3.3)
Combining (3.2) and (3.3), we find that
$\frac{\mathcal{B}^{*}_{n}}{n\mathcal{B}_{n}}\to 1-\rho.$ (3.4)
Hence, in proving Theorem 2, the following is key.
Proposition 7.
We have that $\mathcal{B}^{*}_{n}=2{\mathcal{T}}_{n}$.
We prove Proposition 7 in the next two sections. For now, let us note that our main result follows.
Proof of Theorem 2.
By (1.2) and Proposition 7,
$\frac{\mathcal{B}^{*}_{n}}{4^{n}}\sim\frac{1}{n4^{n}}{2n\choose n}\sim\frac{1}{\sqrt{\pi}n^{3/2}}$
is regularly varying with index $\gamma=-3/2$. Therefore, by (2.3),
$\frac{\mathcal{B}^{*}_{n}}{n\mathcal{B}_{n}}\to\exp(-2\omega).$ (3.5)
Combining (3.4) and (3.5), we find that $\rho=1-\exp(-2\omega)$, as claimed. ∎
4. Combinatorial lemmas
In this section, we prove two combinatorial lemmas, showing that $2{\mathcal{T}}_{n}$ can be described in terms of areas $\alpha$ below lattice paths $L$ and diamond areas $\sigma$ under bridges $B$.
4.1. Lattice paths
Suppose that $L$ is an $\uparrow$,$\rightarrow$ , → lattice path from $(a_{1},b_{1})$ to $(a_{2},b_{2})$, for some integers $a_{1}\leqslant a_{2}$ and $b_{1}\leqslant b_{2}$. We let $\alpha(L)$ denote the area of the region between $L$ and the lines $x=a_{2}$ and $y=b_{1}$.
Definition 8.
We let $\mathcal{N}_{n}$ be the number of lattice paths $L$ from $(0,0)$ to $(n,n)$ such that $\alpha(L)\equiv 0$ mod $n$.
To see the relationship between $\mathcal{N}_{n}$ and ${\mathcal{T}}_{n}$, first note that the area of a lattice path $L$ from $(0,0)$ to $(n,n)$ is
$\alpha(L)=\sum_{i}u_{i},$ (4.1)
summing over the sequence $u_{1}\leqslant\cdots\leqslant u_{n}$, where $u_{i}$ is the number of $\uparrow$ steps before the $i$th $\rightarrow$ step of $L$. In other words, if we picture $L$ as a bar graph, then the $u_{i}$ are the heights of the bars. Hence, $\mathcal{N}_{n}$ is the number of submultisets of $\{0,1,\ldots,n\}$ of size $n$ that sum to 00 mod $n$. On the other hand, recall that ${\mathcal{T}}_{n}$ is equal to the number of submultisets $\{0,1,\ldots,n-1\}$ of size $n$ that sum to 00 mod $n$.
Using these observation, we show the following.
Lemma 9.
We have that $\mathcal{N}_{n}=2{\mathcal{T}}_{n}$.
Proof.
First, we claim that ${\mathcal{T}}_{n}$ is the number of $\uparrow$,$\rightarrow$ , → lattice paths $L$ from $(0,0)$ to $(n,n-1)$ with $\alpha(L)\equiv 0$ mod $n$. Indeed, submultisets $10\leqslant u_{1}\leqslant\cdots\leqslant u_{n}\leqslant n-$ 1 correspond to such $L$ with $\rightarrow$ steps at heights $u_{i}$, and so $\alpha(L)=\sum_{i}u_{i}\equiv 0$ mod $n$.
Let $\mathcal{L}_{n}^{\uparrow}$ (resp. $\mathcal{L}_{n}^{\rightarrow}$) be the number of lattice paths $L$ from $(0,0)$ to $(n,n)$ such that $\alpha(L)\equiv 0$ mod $n$, ending with an $\uparrow$ (resp. $\rightarrow$) step. Then $\mathcal{N}_{n}=\mathcal{L}_{n}^{\uparrow}+\mathcal{L}_{n}^{\rightarrow}$. To conclude, we show that $\mathcal{L}_{n}^{\uparrow}=\mathcal{L}_{n}^{\rightarrow}=\mathcal{T}_{n}$. Indeed, $\mathcal{L}_{n}^{\uparrow}=\mathcal{L}_{n}^{\rightarrow}$ follows by symmetry, reflecting over $x=y$, observing that divisibility by $n$ of the enclosed area is invariant under this reflection. Finally, $\mathcal{L}_{n}^{\uparrow}=\mathcal{T}_{n}$, since if $L$ from $(0,0)$ to $(n,n)$ has last step $\uparrow$, then removing this step yields a corresponding $L^{\prime}$ from $(0,0)$ to $(n,n-1)$ with the same area as $L$. ∎
4.2. Bridges
Next, we connect $\mathcal{N}_{n}$ to bridges.
Definition 10.
We let $\mathcal{N}_{n}^{\prime}$ denote the number of bridges $B$ of length $2n$, with diamond area $\sigma(B)\equiv 0$ mod $n$.
Lemma 11.
We have that $\mathcal{N}_{n}=\mathcal{N}_{n}^{\prime}$.
Proof.
To obtain a correspondence between bridges $B=(B_{0},\ldots,B_{2n})$ of length $2n$ such that $\sigma(B)\equiv 0$ mod $n$ and lattice paths $L$ from $(0,0)$ to $(n,n)$ with $a(L)\equiv 0$ mod $n$, we proceed as follows.
First, let $\Delta_{i}=B_{i}-B_{i-1}$ be the $i$th increment of $B$. For $1\leqslant k\leqslant n$, put
$B^{\mathrm{odd}}_{k}=\sum_{i=1}^{k}\Delta_{2i-1},\quad\quad B^{\mathrm{even}}_{k}=\sum_{i=1}^{k}\Delta_{2i},$
so that $B^{\mathrm{odd}}=(B^{\mathrm{odd}}_{1},\ldots,B^{\mathrm{odd}}_{n})$ and $B^{\mathrm{even}}=(B^{\mathrm{even}}_{1},\ldots,B^{\mathrm{even}}_{n})$ are the walks with the odd and even increments of $B$, respectively.
Next, rotate $B^{\mathrm{odd}}$ counterclockwise by $\pi/4$ to obtain a lattice path $L_{1}$ from $(0,0)$ to $(n-\ell,\ell)$ for some $\ell$ (since $B^{\mathrm{odd}}$ has $n$ steps). Likewise, rotate $B^{\mathrm{even}}$ counterclockwise by $\pi/4$ to obtain a lattice walk $L_{2}$ from $(n-\ell,\ell)$ to $(n,n)$ (since $B$ is a bridge). Let $L$ be the concatenation of $L_{1}$ and $L_{2}$. This procedure is depicted in Figure 4.
Geometric considerations (see Figure 5) imply that
$\alpha(L)=\alpha(L_{1})+\alpha(L_{2})+\ell^{2}.$
Furthermore, observe that, by definition,
$\sigma(B)=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\left(\Delta_{2j-1}+\Delta_{2j}\right)=\frac{1}{2}\sum_{k=1}^{n}(B^{\mathrm{odd}}_{k}+B^{\mathrm{even}}_{k}),$
which equals the signed area enclosed by $B^{\mathrm{odd}}$ and $-B^{\mathrm{even}}$, as depicted in Figure 5. Therefore, it can be seen that
$\sigma(B)=\alpha(L_{1})-[\ell(n-\ell)-\alpha(L_{2})]=\alpha(L)-\ell n.$
Since $\alpha(L)\equiv\sigma(B)$ mod $n$, this completes the proof. ∎
Figure 4. The bijection in Lemma 11.
Figure 5. Calculating areas in Lemma 11.
5. The Lévy–Khintchine transform
We are now in a position to prove our key result Proposition 7, which identifies $\mathcal{B}_{n}^{*}=2{\mathcal{T}}_{n}$ as the Lévy–Khintchine transform of $\mathcal{B}_{n}$.
Proof of Proposition 7.
As discussed, $\mathcal{B}_{n}$ is a renewal sequence, and so $\mathcal{B}_{n}^{*}$ can be described in terms of Lemma 3(1). Combining this with our combinatorial results Lemmas 9 and 11 above, it follows that, to show that $\mathcal{B}_{n}^{*}=2{\mathcal{T}}_{n}$, it suffices to find a bijection:
  • from the set of ordered pairs $(B,i)$, where $B$ is a graphical bridge of length $2n$, whose first irreducible part is of length $2\ell$, and $0\leqslant i<\ell$
  • to the set of bridges $B^{\prime}$ of length $2n$, with diamond area $\sigma(B^{\prime})\equiv 0$ mod $n$.
We claim that such a bijection can be constructed as follows: For each such pair $(B,i)$, let $\phi(B,i)$ be the bridge $B^{\prime}$ obtained from $B$, via a cyclical shift to the right by $2i$. Note that, geometrically, $B^{\prime}$ can be obtained from $B$ by shifting the $x$-axis up or down by some multiple of $2$ and then starting the bridge from a particular intersection point of the new $x$-axis and the bridge $B$. Only the former operation affects the total diamond area. In fact, by (3.1), it can be seen that such a shift changes the area by a multiple of $n$. Since $\sigma(B)=0$, it follows, again by (3.1), that $\sigma(B^{\prime})\equiv 0$ mod $n$. See Figure 6.
Figure 6. A graphical bridge $B$ (top left) of length 10, with first irreducible part of length 8. The bridge $B$ is $\phi(B,0)$. Its other shifts $\phi(B,i)$, for $41\leqslant i<$ 4, are also depicted. All bridges have diamond area divisible by 5. To find the inverse mapping $\phi^{-1}$, we shift the $x$-axis (see dotted line) by some factor of 2 so that the diamond area is equal to 0, and then find the rightmost point (open dot) that starts a graphical sequence. Such a point exists by Raney’s lemma.
On the other hand, for any $B^{\prime}$ with $\sigma(B^{\prime})\equiv 0$ mod $n$, the inverse $\phi^{-1}(B^{\prime})$ is found as follows. Select the unique shift of the $x$-axis for which the diamond area becomes 0, and then choose the rightmost starting point $2n-2i$ for which the resulting bridge $B$ is graphical; then $\phi^{-1}(B^{\prime})=(B,i)$. Such a point exists by Raney’s lemma [17], and by choosing $i$ minimal we ensure that $2i$ is smaller than the length of the first irreducible part of $B$. ∎